OpenStudy (anonymous):

What is the angle of refraction when a ray of light passes from ethanol to air if the incident angle is 34°? The refractive index of ethanol is 1.36 and of air is 1.00.

OpenStudy (ipwnbunnies):

Use Snell's Law. \[n_1 \sin(\Theta_{1}) = n_2 \sin(\Theta_{2})\] Where the left side of the equation is the incident refractive index and angle. The right is the...errr...transmitted info.

OpenStudy (ipwnbunnies):

So, it's going from ethanol to air. n1 = 1.36 n2 = 1.00

OpenStudy (anonymous):

I don't understand any of that ah

OpenStudy (anonymous):

1.36sin(Θ1)=1.00sin(Θ2) but how do i find (Θ1) and (Θ2)

OpenStudy (anonymous):

and like i dunno

OpenStudy (ipwnbunnies):

Θ1 is the incident angle, 34 degrees. \[(1.36)\sin(34 \deg) = (1.00)\sin (\Theta_2)\] I suggest to solve the left side first, use a calculator. Divide by 1, to get sin(Θ2) by itself. Finally, use inverse sine.

OpenStudy (anonymous):

i got 0.719

OpenStudy (ipwnbunnies):

For the left side?

OpenStudy (anonymous):

yeah

OpenStudy (ipwnbunnies):

I got something a little different. o-o

OpenStudy (anonymous):

i don't know how to do reverse sine i have no idea what i am doing

OpenStudy (ipwnbunnies):

I got 0.760. And don't you know how to use a calculator? \[0.760 = \sin(\Theta_2)\] \[\Theta_2 = \sin^{-1}(0.760)\]

OpenStudy (ipwnbunnies):

Do you have your calculator? Press 2nd Key -> SIN. That should be the inverse sine function.

OpenStudy (shannonmcc10):

Answer is 49 degrees